Abstract Submitted for the DNP08 Meeting of The American Physical Society

Sorting Category: 12. (E)

Beam Position Stabilization Using an Active Collimator in Hall D at Jefferson Lab RICHARD JONES, IGOR SENDEROVICH, University of Connecticut, GLUEX COLLABORATION^{1} — The GlueX experiment planned for Hall D at Jefferson Lab relies on the process of coherent bremsstrahlung by 12 GeV electrons in a diamond crystal to produce a secondary beam of 9 GeV photons with a high degree of linear polarization. To achieve optimum polarization, the photon beam must be collimated to a half-angle of 20 μr, which requires a long flight path of 75 m between the radiator and collimator and a means to keep the beam spot centered on the collimator aperture with an accuracy of a few μr. A active collimator has been designed which is capable of monitoring the centroid of the photon beam to within ±200 μm, with a sampling frequency of up to several hundred Hz. A prototype of this this device has been tested in the photon beam in Hall B. Results from this test are presented.

1photon beam working group

Prefer Oral Session X Prefer Poster Session

Richard Jones

richard.t.jones@uconn.edu University of Connecticut

Date submitted: 01 Jul 2008 Electronic form version 1.4